Please help me with this question. Thanks in advance.
An equation and two ordered pairs are given. Show each pair is a solution of the equation, then graph the two pairs to determine another solution.
y = x - 2; (3,1),(- 2,- 4)

Answer by rfer(16322)

y=3-2
y=1
(3,1)
-----------
y=-2-2
y=-4
(-2,-4)